OpenStudy (anonymous):

The cost and revenue functions for a frozen yogurt shop are as follows: C(x)= 400x+400/x+4 and R(x) =100x What is the break even point? If the profit is P(x) then does P (1) and P(4) represent a profit or loss? PLEASE HELP

OpenStudy (dumbcow):

P(x) = R(x) - C(x) P(x) = -300x + 400/x + 4 Break even point is when P(x) = 0 0 = -300x + 400/x + 4 300x^2 = (400+4x) 300x^2 -4x - 400 = 0 divide by 4 75x^2 - x - 100 = 0 use quadratic formula x = (1 +- sqrt(30001)) / 150 x = 1.16, disregard neg answer (can't sell neg amount of product)

OpenStudy (anonymous):

WOW...u r absolutely awesome!!!